The dimensions around a basketball court is 196 feet. If the length of the court is 8 feet more than twice the width, find the dimensions of the court.
Answer by checkley79(3341)

PERIMETER =2L+2W
L=2W+8
196=2(2W+8)+2W
196=4W+16+2W
196=6W+16
6W=196-16
6W=180
W=180/6
W=30 FT WIDE FOR THE WIDTH.
L=2*30+8
L=60+8
L=68 FT. FOR THE LENGTH.
PROOF:
196=2*68+2*30
196=136+60
196=196